OpenStudy (anonymous):

Solve |2x - 5| = 4. {x | x = -4.5 or x = 4.5} {x | x = 0.5 or x = 4.5} {x | 0.5 < x < 4.5}

OpenStudy (anonymous):

second choice is correct

OpenStudy (anonymous):

\[2X-5=\pm4\] 2X=(+4+5)/2 X=0.5 2X=(-4+5)/2 X=4.5